Required information

[The following information applies to the questions displayed below.]

Whitney received $74,800 of taxable income in 2019. All of the income was salary from her employer. What is her income tax liability in each of the following alternative situations? Use Tax Rate Schedule for reference. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

a. She files under the single filing status.

Income tax liability:

b. She files a joint tax return with her spouse. Together their taxable income is $74,800.

Income tax liability:

c. She files as a head of household.

Income tax liability::
Answer #1

(a) She files under the single filing status.

Taxable income = $74,800

Tax payable = [(74,800 - 39,475) x 22%] + 4543 = 7,771.50 + 4543 = $12,314.50

(b) She files a joint tax return with her spouse. Together their taxable income is $74,800.

Taxable income = $74,800

Tax payable = [(74,800 – 19,400) x 12%] + 1940 = 6,648 + 1940 = $8,588

(c) She files as a head of household

Taxable income = $74,800

Tax payable = [(74,800 – 52,850) x 22%] + 6065 = 4,829 + 6065 = $10,894